Given the following mixture of two compounds 50.00 mL of X (MW =75.00 g/mol)(density 1.257 g/mL) and 740.00 mL of Y (51.00 g/mol))(density 1.002 g/mL), calculate the mole fraction of X.
1
Expert's answer
2020-08-31T07:04:20-0400

Mole fraction of X is 0.0545.


m (X) =50.00mL ×1.257g/mL=62.85 g


n (X) =62.85 g/75g/mol= 0.838 mol


m (Y) = 740 mL × 1.002g/mL=741.48 g


n (Y)=741.48 g/51g/mol= 14.54 mol


Mole fraction (X)=0.838/(0.838+14.54)=0.0545